OpenStudy (anonymous):

The world population double every 30 years, and population on Jan 1, 1970 is 5 billion. In what year will population reach 12 billion?

OpenStudy (anonymous):

solve \[12=5\times 2^{\frac{t}{30}}\] for t

OpenStudy (anonymous):

divide by 5 get \[2.4=2^{\frac{t}{30}}\] write in equivalent logartihmic form get \[\frac{t}{30}=\frac{\ln(2.4)}{\ln(2)}\] then multiply by 30 get \[t=\frac{30\ln(2.4)}{\ln(2)}\] then use a calculator